OpenStudy (ebydaqueen19):

SOMEONE PLZ HELP ME..BIG TEST!!! At 27C and 750 torr, both a 16 g sample of methane gas and 16 g sample of oxygen gas will have the same: A. average molecular velocities B. average molecular kinetic energies C. number of gaseous particles D. total gaseous volumes E. average effusion rates

OpenStudy (kc_kennylau):

C, D and E are related to moles, so they must be wrong

OpenStudy (ebydaqueen19):

it says b is the answer

OpenStudy (ebydaqueen19):

but i want to know why

OpenStudy (ebydaqueen19):

can u explain a little bit plz

OpenStudy (kc_kennylau):

Average molecular kinetic energy is related to the temperature

OpenStudy (kc_kennylau):

Actually average molecular kinetic energy IS temperature

OpenStudy (ebydaqueen19):

yeah tht is right...

OpenStudy (ebydaqueen19):

but they have the same pressure as well

OpenStudy (ebydaqueen19):

why is tht

OpenStudy (kc_kennylau):

Since they have the same pressure

OpenStudy (kc_kennylau):

The bag containing the methane gas will be smaller in size

OpenStudy (ebydaqueen19):

so thts why it couldnt be same gas particles or same volume.. right

OpenStudy (kc_kennylau):

Yep

OpenStudy (ebydaqueen19):

and effusion rate is far too wild of an answer right

OpenStudy (ebydaqueen19):

effusion rate has nothing to do with it, huh?

OpenStudy (kc_kennylau):

Effusion rate is proportional to the number of moles

OpenStudy (kc_kennylau):

I mean the square root of the ratio

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Effusion

OpenStudy (kc_kennylau):

\[\frac{Rate_1}{Rate_2}=\sqrt{\frac{M_1}{M_2}}\]

OpenStudy (ebydaqueen19):

oh ok thx

OpenStudy (kc_kennylau):

no problem

OpenStudy (ebydaqueen19):

would u mind helping on a few more problems just a few

OpenStudy (ebydaqueen19):

had to pull an all nighter...lol u noe how tht is

OpenStudy (kc_kennylau):

okay open another problem

OpenStudy (ebydaqueen19):

altrite thx